OpenStudy (helpmeplzz):

what is the 1,2,6,24,120,720... nth term sequence

OpenStudy (anonymous):

do you want to learn the material or just get the answer

OpenStudy (anonymous):

t1=1 t2=t1*2=2 t3=t2*3=6 t4=t3*4=24 t5=t4*5=120

OpenStudy (anonymous):

Do u see a patern

OpenStudy (anonymous):

t6=t5*6=720 t7=t6*7=720*7=

OpenStudy (helpmeplzz):

I want to know the n th term formula

OpenStudy (anonymous):



OpenStudy (helpmeplzz):

it does not work..

OpenStudy (helpmeplzz):

@sauravshakya